# Magneto-Thomson and transverse Thomson effects in an interacting hadron gas in the presence of an external magnetic field

### v1

@@ Abstract
The universality of electric charge as a quantum number allows thermoelectric properties to manifest across diverse systems, starting from a hot quantum chromodynamics matter in heavy-ion collisions at a high energy scale to semiconductors in condensed matter systems at a low energy scale. In this work, we explore the emergence of magneto-transport phenomena, specifically the magneto-Thomson and transverse Thomson effects, in a hot and dense hadronic medium produced in relativistic heavy-ion collisions at Relativistic Heavy Ion Collider and Large Hadron Collider energies. These phenomena arise due to the combined influence of temperature gradients and non-zero baryon chemical potential, particularly in the presence of an external magnetic field. Using the relativistic Boltzmann transport equation within the relaxation time approximation, we analyze the behavior of the hadronic medium considering different frameworks of hadron resonance gas models. The presence of external magnetic fields breaks the isotropy of the thermoelectric transport coefficient matrix, giving rise to new components of the Thomson coefficient, namely, magneto-Thomson and transverse Thomson coefficients. For the first time, we estimate the magneto-Thomson and transverse Thomson coefficients, which originate from the temperature dependence of the magneto-Seebeck coefficient and Nernst coefficient, respectively, in hadron gas under the influence of a magnetic field. Our findings provide a novel perspective on the higher-order thermoelectric properties of the hot and dense hadronic medium in the context of heavy-ion collisions.
@@ I Introduction
The dynamics of thermoelctricity is affected by the presence of a nonzero magnetic field because of the experience of the Lorentz force by charge carriers. Peripheral heavy-ion collisions are capable of generating extremely strong magnetic fields as a result of the relativistic motion of spectator protons. In such events, the magnetic field strength can peak around $m_{\pi}^{2}\sim 10^{18}$ G for Au-Au collisions at RHIC and reach up to $15~{}m_{\pi}^{2}$ in Pb-Pb collisions at the LHC [22, 23]. These fields tend to dissipate almost instantly in a vacuum, but the situation changes when a thermalized conducting medium, such as the quark-gluon plasma, is present. Due to its finite electrical conductivity, the medium supports the induction of secondary magnetic fields aligned with the initial external field, as described by Faraday’s law [24]. This results in a slower decay of the magnetic field, allowing it to play a significant role in influencing the complete evolution of the fireball, particularly its thermodynamic and transport properties. The influence of magnetic fields on the conductivity of QCD medium, both electrical and thermal, has been extensively addressed in the literature [4, 25, 26]. Under such magnetic field conditions, leading-order thermoelectric transport becomes directionally dependent, giving rise to the magneto-Seebeck and Nernst coefficients. Unlike the Seebeck coefficient, which responds to longitudinal temperature gradients alone, the Nernst coefficient is a transverse response that emerges only in the presence of magnetic fields. It quantifies the generation of an electric current perpendicular to both the thermal gradient and the magnetic field [12]. Previous studies have analyzed this behavior for the hadronic medium using hadron resonance gas models [14, 16], while the study [27] explores the impact of magneto-thermoelectric effects on conserved charge diffusion. The thermoelectric behavior of the QGP has also been explored in both isotropic [28] and anisotropic [29, 30] media, largely within perturbative QCD formalisms, which are valid at very high temperatures. The temperature dependence of the leading-order thermoelectric coefficients in a magnetic field gives rise to the higher-order thermoelectric coefficients. In our previous study [13], we calculated the Thomson coefficient of interacting hot hadron gas, which originates due to the temperature dependence of the Seebeck coefficient. The Thomson effect can be understood as originating from the simultaneous operation of the Seebeck and Peltier effects [31], and appears when the medium has a temperature dependence of the Seebeck coefficient. If the charge current flows through the spatial gradient of the Seebeck coefficient, the self-induced Peltier effect modulates the temperature in response to the charge current [32]. In the current study, we calculate the Thomson effect for hardon gas in the presence of a magnetic field. The presence of a magnetic field creates an anisotropy in the medium and introduces magneto-Thomson and transverse Thomson coefficients. For the first time, we calculate the magneto-Thomson coefficient and transverse Thomson coefficient for the hot and dense hadron gas. The transverse Thomson effect originates due to the presence of the Nernst effect in the presence of a magnetic field [33]. In the absence of a magnetic field, the magneto-Thomson coefficient reduces to the Thomson coefficient, but the transverse Thomson coefficient vanishes completely.
@@ II.2 Leading-order thermoelectric coefficients
and $\nabla^{\mu}=\Delta^{\mu\nu}\partial_{\nu}$ is the spatial gradient orthogonal to $u^{\mu}$, with $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$. Here, $\epsilon^{ijk}$ denotes the Levi-Civita symbol, which is used to represent antisymmetric tensor components.
@@
$\vec{v}_{i}.\bigg{[}-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}\bigg{(}\frac{\omega_{i}-\beta_{i}h}{T}\bigg{)}\vec{\nabla}T\bigg{]}+q_{i}(\vec{E}.\vec{v}_{i})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}$
$-q_{i}\vec{v}_{i}.(\vec{\Omega}\times\vec{B})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}=-\frac{\omega_{i}}{\tau^{i}_{R}}(\vec{v}_{i}.\vec{\Omega})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (41)
Where $\vec{v}_{i}=\vec{k}_{i}/\omega_{i}$ is the relativistic particle velocity. By substituting $\vec{\Omega}$ from Eq.(40) into Eq.(41), we can expand it analogously to the nonrelativistic case as,
$q_{i}(\vec{E}.\vec{v}_{i})-\alpha_{1}q_{i}\vec{v}_{i}.(\vec{E}\times\vec{B})-\alpha_{3}q_{i}(\vec{E}.\vec{B})(\vec{v}_{i}.\vec{B})+\alpha_{3}$
$q_{i}(\vec{v}_{i}.\vec{E})-\alpha_{4}q_{i}\vec{v}_{i}.(\vec{\nabla}T\times\vec{B})-\alpha_{5}q_{i}(\vec{\nabla}T.\vec{B})(\vec{v}_{i}.\vec{B})$
$+\alpha_{5}q_{i}(\vec{v}_{i}.\vec{\nabla}T)-\alpha_{6}q_{i}(\vec{\nabla}T.\vec{B})(\vec{v}_{i}.\vec{E})+\alpha_{6}q_{i}(\vec{E}.\vec{B})(\vec{v}_{i}.\vec{\nabla}T)$
$-\bigg{(}\frac{\omega_{i}-\beta_{i}h}{T}\bigg{)}(\vec{v}_{i}.\vec{\nabla}T)=-\frac{\omega_{i}}{\tau^{i}_{R}}\bigg{[}\alpha_{1}(\vec{v}_{i}.\vec{E})+\alpha_{2}(\vec{v}_{i}.\vec{B})$
$+\alpha_{3}\vec{v}_{i}.(\vec{E}\times\vec{B})+\alpha_{4}(\vec{v}_{i}.\vec{\nabla}T)+\alpha_{5}\vec{v}_{i}.(\vec{\nabla}T\times\vec{B})+$
$\alpha_{6}\vec{v}_{i}.(\vec{\nabla}T\times\vec{E})\bigg{]}.$ (42)
@@
$\alpha_{2}=\alpha_{3}\tau^{i}_{R}\Omega_{c_{i}}(\vec{E}.\vec{B})+b\tau^{i}_{R}\Omega_{c_{i}}(\vec{\nabla}T.\vec{B}),$ (44)
@@
$\delta f_{i}$ $=\frac{\tau^{i}_{R}}{1+(\Omega_{c_{i}}\tau^{i}_{R})^{2}}\bigg{[}q_{i}\bigg{\{}(\vec{v}_{i}.\vec{E})+(\Omega_{c_{i}}\tau^{i}_{R})\vec{v}_{i}.(\vec{E}\times\vec{B})+$
$(\Omega_{c_{i}}\tau^{i}_{R})^{2}(\vec{E}.\vec{B})(\vec{v}_{i}.\vec{B})\bigg{\}}-\bigg{(}\frac{\omega_{i}-\beta_{i}h}{T}\bigg{)}\bigg{\{}(\vec{v}_{i}.\vec{\nabla}T)+$
$(\Omega_{c_{i}}\tau^{i}_{R})\vec{v}_{i}.(\vec{\nabla}T\times B)+(\Omega_{c_{i}}\tau^{i}_{R})^{2}(\vec{\nabla}T.\vec{B})(\vec{v}_{i}.\vec{B})\bigg{]}(-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}).$ (51)
@@
Here, to simplify further calculations, we can choose the magnetic field along the $z$ direction. The direction of the electric field and the temperature gradient are perpendicular to the $z$ axis, i.e., it is in the $x-y$ plane. Under these conditions, the components of the electric current in the $x-y$ plane are given as,
@@
$(-)\frac{\partial f_{i}^{0}}{\partial\omega_{i}}-\sum_{i}\frac{g_{i}q_{i}}{3T}\int\frac{d^{3}k_{i}}{(2\pi)^{3}}\frac{v_{i}^{2}\tau^{i}_{R}\left(\omega_{i}-\beta_{i}h\right)}{1+(\Omega_{c_{i}}\tau)^{2}}\bigg{]}$
@@
Here, we have identified the ohmic-like component of electrical conductivity in the presence of a magnetic field and the Hall-like component of electrical conductivity as $\sigma_{el}=\sum_{i}q_{i}^{2}H_{1_{i}}$ and $\sigma_{H}=\sum_{i}q_{i}^{2}H_{2_{i}}$ respectively [12]. The integrals $\mathcal{I}_{31}$ and $\mathcal{I}_{42}$ in Eqs.(65) and (66) are defined as $\mathcal{I}_{31}=\sum_{i}q_{i}(H_{3_{i}}-\beta_{i}hH_{1_{i}})$ and $\mathcal{I}_{42}\equiv\sum_{i}q_{i}(H_{4_{i}}-\beta_{i}hH_{2_{i}})$. Note that in the absence of a magnetic field, integrals $H_{2_{i}}$ and $H_{4_{i}}$ are identically zero. Hence, the normalized Nernst coefficient vanishes in the absence of a magnetic field, and the magneto-Seebeck coefficient reduces to the Seebeck coefficient in the absence of a magnetic field [13]. The coefficients $S_{B}$ and $NB$ have temperature dependence as shown in Eqs. 65 and 66, hence the higher-order thermoelectric coefficients are introduced for the medium.
@@ II.3 Higher-order thermoelectric coefficients
In the above calculation, for the sake of simplicity, we take the thermal averaged relaxation time after integrating the energy-dependent relaxation time over the equilibrium distribution function. The thermal averaged relaxation time ($\tau_{R}^{i}$) for the ith hadron species can be expressed in terms of scattering cross-section as [16],
@@
$\langle\sigma_{ij}v_{ij}\rangle$ $=\frac{\sigma}{8Tm_{i}^{2}m_{j}^{2}\mathcal{K}_{2}(m_{i}/T)\mathcal{K}_{2}(m_{j}/T)}~{}\times$
$\int_{(m_{i}+m_{j})^{2}}^{\infty}ds~{}\times$

### v2

@@ Abstract
The universality of electric charge as a quantum number allows thermoelectric properties to manifest across diverse systems, starting from a hot quantum chromodynamic matter in heavy-ion collisions at a high energy scale to semiconductors in condensed matter systems at a low energy scale. In this work, we explore the emergence of magneto-transport phenomena, specifically the magneto-Thomson and transverse Thomson effects, in a hot and dense hadronic medium produced in relativistic heavy-ion collisions at the Relativistic Heavy Ion Collider and Large Hadron Collider energies. These phenomena arise due to the combined influence of temperature gradients and non-zero baryon chemical potential, particularly in the presence of an external magnetic field. Using the relativistic Boltzmann transport equation within the relaxation time approximation, we analyze the behavior of the hadronic medium considering different frameworks of hadron resonance gas models. The presence of external magnetic fields breaks the isotropy of the thermoelectric transport coefficient matrix, giving rise to new components of the Thomson coefficient, namely, magneto-Thomson and transverse Thomson coefficients. For the first time, we estimate the magneto-Thomson and transverse Thomson coefficients, which originate from the temperature dependence of the magneto-Seebeck coefficient and Nernst coefficient, respectively, in hadron gas under the influence of a static and time-varying magnetic field. Our findings provide a novel perspective on the higher-order thermoelectric properties of the hot and dense hadronic medium in the context of heavy-ion collisions.
@@ I Introduction
The dynamics of thermoelctricity is affected by the presence of a nonzero magnetic field because of the experience of the Lorentz force by charge carriers. Peripheral heavy-ion collisions are capable of generating extremely strong magnetic fields as a result of the relativistic motion of spectator protons. In such events, the magnetic field strength can peak around $m_{\pi}^{2}\sim 10^{18}$ G for Au-Au collisions at RHIC and reach up to $15~m_{\pi}^{2}$ in Pb-Pb collisions at the LHC [22, 23]. These fields tend to dissipate almost instantly in a vacuum, but the situation changes when a thermalized conducting medium, such as the quark-gluon plasma, is present. Due to its finite electrical conductivity, the medium supports the induction of secondary magnetic fields aligned with the initial external field, as described by Faraday’s law [24]. This results in a slower decay of the magnetic field, allowing it to play a significant role in influencing the complete evolution of the fireball, particularly its thermodynamic and transport properties. In the hadronic phase, implementing a time-varying magnetic field requires an evolving description of the medium, which is not inherently present in the static thermodynamic framework of the HRG model. Unlike the QGP phase, where hydrodynamic profiles can be incorporated to describe the space–time evolution and hence the decay of the magnetic field, the HRG model in its conventional form lacks such dynamic input. However, in the hadronic phase, the space-time evolution of the matter and hence of the electromagnetic fields can be incorporated by means of kinetic theory. Therefore, we adopt both static and dynamical magnetic field pictures in the HRG phase. We use two representative strengths for the static picture ($eB=0.1$ and $1.0~m_{\pi}^{2}$[25] and the dynamical picture with magnetic decay parameter $\tau_{B}$ = 6 fm, to qualitatively mimic the possible range of field strengths during the late stages of the collision. The influence of magnetic fields on the conductivity of QCD medium, both electrical and thermal, has been extensively addressed in the literature [4, 26, 27]. Under such magnetic field conditions, leading-order thermoelectric transport becomes directionally dependent, giving rise to the magneto-Seebeck and Nernst coefficients. Unlike the Seebeck coefficient, which responds to longitudinal temperature gradients alone, the Nernst coefficient is a transverse response that emerges only in the presence of magnetic fields. It quantifies the generation of an electric current perpendicular to both the thermal gradient and the magnetic field [12]. Previous studies have analyzed this behavior for the hadronic medium using hadron resonance gas models [14, 16], while the study [28] explores the impact of magneto-thermoelectric effects on conserved charge diffusion. The thermoelectric behavior of the QGP has also been explored in both isotropic [29] and anisotropic [30, 31] media, largely within perturbative QCD formalisms, which are valid at very high temperatures. The temperature dependence of the leading-order thermoelectric coefficients in a magnetic field gives rise to the higher-order thermoelectric coefficients. In our previous study [13], we calculated the Thomson coefficient of interacting hot hadron gas, which originates due to the temperature dependence of the Seebeck coefficient. The Thomson effect can be understood as originating from the simultaneous operation of the Seebeck and Peltier effects [32], and appears when the medium has a temperature dependence of the Seebeck coefficient. If the charge current flows through the spatial gradient of the Seebeck coefficient, the self-induced Peltier effect modulates the temperature in response to the charge current [33]. In the current study, we calculate the Thomson effect for hardon gas in the presence of a magnetic field. The presence of a magnetic field creates an anisotropy in the medium and introduces magneto-Thomson and transverse Thomson coefficients. For the first time, we calculate the magneto-Thomson coefficient and transverse Thomson coefficient for the hot and dense hadron gas. The transverse Thomson effect originates due to the presence of the Nernst effect in the presence of a magnetic field [34]. In the absence of a magnetic field, the magneto-Thomson coefficient reduces to the Thomson coefficient, but the transverse Thomson coefficient vanishes completely. In principle, strong magnetic fields, such as those generated in the early stages of non-central heavy-ion collisions at LHC energies, can modify hadronic properties through mechanisms such as Landau quantization, which may in turn influence hadron-hadron interactions. It should be noted that, besides Landau quantization, strong magnetic fields can also modify hadron–hadron interactions by affecting the interaction vertices and propagators, particularly for charged hadrons [35, 36, 37, 38]. Such modifications could in principle influence scattering cross-sections and relaxation times, thereby altering transport coefficients. A detailed quantum-field-theoretic treatment of these effects lies beyond the scope of the present kinetic-theory framework. Here, the effect of the magnetic field enters into the transport equations through the cyclotron frequency of individual hadrons, while the hadronic scattering cross-sections are taken to be field-independent within the hard‑sphere approximation. The effects arising from Landau quantization for the charged hadrons are also explicitly considered in this study.
@@ II.2 Leading-order thermoelectric coefficients
and $\nabla^{\mu}=\Delta^{\mu\nu}\partial_{\nu}$ is the spatial gradient orthogonal to $u^{\mu}$, with $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$. Here, $\epsilon^{\mu\nu\rho\sigma}$ denotes the Levi-Civita symbol, which is used to represent antisymmetric tensor components.
@@
$\vec{v}_{i}\cdot\bigg[-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}\bigg(\frac{\omega_{i}-\beta_{i}h}{T}\bigg)\vec{\nabla}T\bigg]+q_{i}(\vec{E}\cdot\vec{v}_{i})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}$
$-q_{i}\vec{v}_{i}\cdot(\vec{\Omega}\times\vec{B})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}=-\frac{\omega_{i}}{\tau^{i}_{R}}(\vec{v}_{i}\cdot\vec{\Omega})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (41)
Where $\vec{v}_{i}=\vec{k}_{i}/\omega_{i}$ is the relativistic particle velocity and $h=\frac{\epsilon+P}{n}$ is the enthalpy per baryon, $\epsilon,P$, and $n$ are total energy density, total pressure, and net baryon density of the system, respectively. Here, the enthalpy per baryon is induced in the picture following the Gibbs-Duhem relation. Using the Gibbs–Duhem relation, the gradients of the inverse temperature and the baryon chemical potential are combined, allowing the gradient of the equilibrium distribution function to be rewritten solely in terms of the temperature gradient, with the enthalpy per baryon naturally emerging as the relevant thermodynamic quantity [5, 12, 16, 65].
By substituting $\vec{\Omega}$ from Eq.(40) into Eq.(41), we can expand it analogously to the nonrelativistic case as,
$q_{i}(\vec{E}\cdot\vec{v}_{i})-\alpha_{1}q_{i}\vec{v}_{i}\cdot(\vec{E}\times\vec{B})-\alpha_{3}q_{i}(\vec{E}\cdot\vec{B})(\vec{v}_{i}\cdot\vec{B})+\alpha_{3}$
$q_{i}(\vec{v}_{i}\cdot\vec{E})-\alpha_{4}q_{i}\vec{v}_{i}\cdot(\vec{\nabla}T\times\vec{B})-\alpha_{5}q_{i}(\vec{\nabla}T\cdot\vec{B})(\vec{v}_{i}\cdot\vec{B})+\alpha_{5}$
$q_{i}(\vec{v}_{i}\cdot\vec{\nabla}T)-\alpha_{6}q_{i}(\vec{\nabla}T\cdot\vec{B})(\vec{v}_{i}\cdot\vec{E})+\alpha_{6}q_{i}(\vec{E}\cdot\vec{B})(\vec{v}_{i}\cdot\vec{\nabla}T)$
$-\bigg(\frac{\omega_{i}-\beta_{i}h}{T}\bigg)(\vec{v}_{i}\cdot\vec{\nabla}T)=-\frac{\omega_{i}}{\tau^{i}_{R}}\bigg[\alpha_{1}(\vec{v}_{i}\cdot\vec{E})+\alpha_{2}(\vec{v}_{i}\cdot\vec{B})$
$+\alpha_{3}\vec{v}_{i}\cdot(\vec{E}\times\vec{B})+\alpha_{4}(\vec{v}_{i}\cdot\vec{\nabla}T)+\alpha_{5}\vec{v}_{i}\cdot(\vec{\nabla}T\times\vec{B})+$
$\alpha_{6}\vec{v}_{i}\cdot(\vec{\nabla}T\times\vec{E})\bigg].$ (42)
@@
$\alpha_{2}=\alpha_{3}\tau^{i}_{R}\Omega_{c_{i}}(\vec{E}\cdot\vec{B})+b\tau^{i}_{R}\Omega_{c_{i}}(\vec{\nabla}T\cdot\vec{B}),$ (44)
@@
$\delta f_{i}$ $=\frac{\tau^{i}_{R}}{1+(\Omega_{c_{i}}\tau^{i}_{R})^{2}}\bigg[q_{i}\bigg\{(\vec{v}_{i}\cdot\vec{E})+(\Omega_{c_{i}}\tau^{i}_{R})\vec{v}_{i}\cdot(\vec{E}\times\vec{B})+$
$(\Omega_{c_{i}}\tau^{i}_{R})^{2}(\vec{E}\cdot\vec{B})(\vec{v}_{i}\cdot\vec{B})\bigg\}-\bigg(\frac{\omega_{i}-\beta_{i}h}{T}\bigg)\bigg\{(\vec{v}_{i}\cdot\vec{\nabla}T)+$
$(\Omega_{c_{i}}\tau^{i}_{R})\vec{v}_{i}\cdot(\vec{\nabla}T\times B)+(\Omega_{c_{i}}\tau^{i}_{R})^{2}(\vec{\nabla}T\cdot\vec{B})(\vec{v}_{i}\cdot\vec{B})\bigg\}\bigg](-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}).$ (51)
@@
Here, to simplify further calculations, we can choose the magnetic field along the $z$ direction. The direction of the electric field and the temperature gradient are perpendicular to the $z$ axis, i.e., it is in the xyx-y plane. Under these conditions, the components of the electric current in the xyx-y plane are given as,
@@
Here, we have identified the ohmic-like component of electrical conductivity in the presence of a magnetic field and the Hall-like component of electrical conductivity as $\sigma_{el}=\sum_{i}q_{i}^{2}H_{1_{i}}$ and $\sigma_{H}=\sum_{i}q_{i}^{2}H_{2_{i}}$ respectively [12]. The integrals $\mathcal{I}_{31}$ and $\mathcal{I}_{42}$ in Eqs.(65) and (66) are defined as $\mathcal{I}_{31}=\sum_{i}q_{i}(H_{3_{i}}-\beta_{i}hH_{1_{i}})$ and $\mathcal{I}_{42}=\sum_{i}q_{i}(H_{4_{i}}-\beta_{i}hH_{2_{i}})$. Note that in the absence of a magnetic field, integrals $H_{2_{i}}$ and $H_{4_{i}}$ are identically zero. Hence, the normalized Nernst coefficient vanishes in the absence of a magnetic field, and the magneto-Seebeck coefficient reduces to the Seebeck coefficient in the absence of a magnetic field [13]. The coefficients $S_{B}$ and $NB$ have temperature dependence as shown in Eqs. 65 and 66, hence the higher-order thermoelectric coefficients are introduced for the medium.
@@ II.3 Higher-order thermoelectric coefficients
In the above calculation, for the sake of simplicity, we take the thermal-averaged relaxation time after integrating the energy-dependent relaxation time over the equilibrium distribution function. The thermal averaged relaxation time ($\tau_{R}^{i}$) for the ith hadron species can be expressed in terms of scattering cross-section as [16],
@@
$\langle\sigma_{ij}v_{ij}\rangle$ $=\frac{\sigma}{8Tm_{i}^{2}m_{j}^{2}\mathcal{K}_{2}(m_{i}/T)\mathcal{K}_{2}(m_{j}/T)}$$\times$
$\int_{(m_{i}+m_{j})^{2}}^{\infty}ds$$\times$